6x^2 + 3x - 18 = 0
1
Expert's answer
2011-08-13T08:01:24-0400
6x2 + 3x - 18 = 0
D = 9 + 4*9*18 = 441 = 212
x1 = (-3 +21)/(2*6) = 18/ 12 = 3/2
x2 = (-3 - 21)/(2*6) = -24/12 = 2